OpenStudy (anonymous):

5/2x+15/2=15 x = −3 x = −1.5 x = 0 x = 3

OpenStudy (goformit100):

x = 3 IS THE answer you got ?

OpenStudy (campbell_st):

subtract 15/2 from both sides then you have 5/2 x = 15/2 multiply both sides by 2 then divide both sides by 5 to find x

OpenStudy (anonymous):

hey cn u help and its not 3

OpenStudy (mathstudent55):

(5/2)x+15/2=15 (5/2)x = 15 - 15/2 (5/2)x = 15/2 This much campbell_st already showed you Multiply both sides by 2 5x = 15 Divide both sides by 5 x = 3

OpenStudy (anonymous):

the keeps saying its wrong but thanks

OpenStudy (mathstudent55):



OpenStudy (anonymous):

yes

OpenStudy (anonymous):

i put -3 thats y but can u help with this

OpenStudy (mathstudent55):